Given the function f(x)=4 x^2 + x - 2. Calculate the following values:
f(0)=
f(2)=
f(-2)=
f(x+1)=
f(-x)=

Answer by checkley77(12844)

f(x)=4 x^2 + x - 2. Calculate the following values:
f(0)=4*0^2+0-2=-2
f(2)=4*2^2+2-2=4*4=16
f(-2)=4(-2)^2-2-2=4*4-4=16-4=12
f(x+1)=4(x+1)^2+(x+1)-2=4(x^2+2x+1)+x+1-2=4x^2+8x+4+x-1=4x^2+9x+3
f(-x)=4(-x)^2-x-2=4x^2-x-2